It has long been stated that the mean temperature of humans is 98.6°F. However, two researchers currently involved in the subject thought that the mean temperature of humans is less than 98.6°F. They measured the temperatures of 56 healthy adults 1 to 4 times daily for 3 days, obtaining 250 measurements. The sample data resulted in a sample mean of 98.3°F and a sample standard deviation of 1°F. Use the P-value approach to conduct a hypothesis test to judge whether the mean temperature of humans is less than 98.6°F at the α=0.01 level of significance. 

State the hypotheses. 
Answer #1

Ho :   µ =   98.6      
Ha :   µ <   98.6      
              
Level of Significance ,    α =    0.01      
sample std dev ,    s =    1      
Sample Size ,   n =    250      
Sample Mean,    x̅ =   98.3      
              
degree of freedom=   DF=n-1=   249      
              
Standard Error , SE =   s/√n =   0.0632      
              
t-test statistic=   (x̅ - µ )/SE =    -4.7434      
              
  
p-Value   =   0.0000 (using student t table)
Conclusion:     p-value<α, Reject null hypothesis           
              
there is enough evidence to support the claim of researcher